Mark is 2/5 the age of his brother. if his brother is 25 years old , which equation could be use to determine m , Mark's age?
Answer by stanbon(75887)

Mark is 2/5 the age of his brother. if his brother is 25 years old , which equation could be use to determine m , Mark's age?
---------
Equations:
m = (2/5)b
b = 25
-----
Mark = (2/5)25 = 10
----------------------
Cheers,
Stan H.